Minutes of the 11th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer),

Absent(with apologies) Conor McDermottroe (Public Relations Officer), Absent Alan Smyth (Chief Range Officer),

Jason Sibley,

Deirdre Ninaber (Ladies’ Coordinator) Date 11th October 2001 1900 – 2030 Distribution DURC Committee

Overview of topics

1 Committee Meetings 1 2 Jason 1 3 Treasurers Report 2 4 Club size 2 5 Email mailing lists 2 6 A&E report 2 7 Secretary’s Report 2 8 NonCapital Grant 2

1 Committee Meetings

1 Committee Meetings

Steven suggested fortnightly meetings to keep the emphasis of the committee members’ work on their shooting as opposed to administration. Ben noted that we often needed weekly meetings, Mark pointed out that the constitution required them.

2 Jason

Should Jason be voted off committee due to his prolonged absence? Steven pointed out the stress on Jason due to super-supplementals and agreed to ask Jason if he wanted to remain on committee. Task: Ask Jason about remaining on Committee (Steven)

3 Treasurers Report

Ben reported on the results from the NonCapital grant meeting :

Change Category  Amount Granted
Up £50 Non-cap equipment 2350 euros
Up £5 Travel  381 euros
Up £60 Fees  1100 euros
Coaching  222 euros

Hire, etc. was not applied for. Mark asked where the National Squad sessions came from, Ben replied Coaching.

John pointed out that we had more members than last year, the same amount of money, and that this amount wasn’t enough last year. Everyone agreed belt-tightening was necessary. All were unhappy with the level of funding.

4 Club size

John suggested that we drive to get more members to be offically the largest sports club in college for the extra political clout and possible sponsorship. Ben reports that DUEC will bring at least 12 members, and reported on their request for training. John stresses that logistical planning was essential to prevent a repeat of last year’s fiasco.

5 Email mailing lists

The mailing lists were discussed, no conclusion was reached as to which to use.

6 A&E report

Pellets are needed. A tube has been ordered from Dave Cooney. Mark asked if we were dropping the idea of alternative suppliers. Consensus was that we had to until funds were raised enough to allow us to pursue that option again. £3000 is needed for that.

7 Secretary’s Report

7 Secretary’s Report

Mark noted the article in the Evening Herald on the WTSC Open which he had not managed to secure a copy of; The CARs written for the WTSC open; and reported on the new security arrangements in EARC.

8 NonCapital Grant

A long discussion was held on the prioritisation of the equipment on the noncapital grant application. Priority 1 items came to 1100 euros, Priority 2 came to 445 euros and Priority 3 to 659 euros, leaving a surpleus of only a few hundred euros. This was quite tight. Mark suggested appealing to DUCAC’s treasurer, Ben and John held out little hope. Sponsorship was discussed, including the possiblility of Eley sponsorship in the form of ammo. It was decided however, to not place the orders for our noncap equipment until later in the year both to try to get a better price by making one order for Capital and NonCapital items at once and to do so direct to the manufacturer. A new list of Priority 0 items was drawn up, consisting mainly of kit bags. John to price and locate these items. Task: Source and price the items on the priority zero list (John)

Mark wanted to add torque wrenches to the priority zero list. The current ones were tested and found to be adaquate for our needs.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

8 NonCapital Grant

Tasks agreed on at meeting

Steven Ask Jason about remaining on Committee (11th October 2001) [/141] /2 John Source and price the items on the priority zero list (11th October

2001) [/205] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2